How many moles of carbon are in 3.08*10^23 molecules of C6H12O6
1
Expert's answer
2019-03-05T06:19:22-0500

n(C6H12O6) = N(C6H12O6)/NA = 3.08*10^23 / 6.02*10^23 mol-1 ≈ 0.5116 mol.

n(C) = 6*n(C6H12O6) = 6*0.5116 mol = 3.0696 mol ≈ 3.07 mol.